OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). ?? how do u do this?

OpenStudy (karatechopper):

well what is your slope?

OpenStudy (anonymous):

Karate!!

OpenStudy (karatechopper):

GT!!

OpenStudy (karatechopper):

anyways jazlin what is your slope?

OpenStudy (anonymous):

hpw would i find the slope?

OpenStudy (anonymous):

To find slope, you need to re-arrange the equation into y=mx+b form where m is the slope.

OpenStudy (anonymous):

i did that but i dont know how to completely slove it like this is what i put -2=m(4)+6

OpenStudy (anonymous):

sorry i ment +b

OpenStudy (anonymous):

3x-y=-3 => y = 3x+3 So, slope = 3.

OpenStudy (anonymous):

Now, you need to find the equation of the line parallel to this line and passing through point (4,-2).

OpenStudy (anonymous):

Parallel lines have same slope. So, the new line must have a slope of 3. So, it must have the equation: y = 3x + b. To find b, you should use the fact that this line passes through point (4,-2).

OpenStudy (anonymous):

so the it would be -2=3(4)+b nd i have to solve it ?

OpenStudy (anonymous):

does b=-6?

OpenStudy (anonymous):

You got part of it. -2 = 3*4+b => -2 - 12 = b = -14 So, new line is: y = 3x-14

OpenStudy (anonymous):

ohh thanks